HBM, Inc has the following capital structure:

Assets $ 450,000 Debt $ 112,500
Preferred stock 45,000
Common stock 292,500

The common stock is currently selling for $15 a share, pays a cash dividend of $0.90 per share, and is growing annually at 3 percent. The preferred stock pays a $10 cash dividend and currently sells for $96 a share. The debt pays interest of 7.0 percent annually, and the firm is in the 30 percent marginal tax bracket.

a) What is the after-tax cost of debt? Round your answer to two decimal places.

_____ %

b) What is the cost of preferred stock? Round your answer to two decimal places.

_______%

c) What is the cost of common stock? Assume that the current $0.90 dividend grows by 3 percent during the year. Round your answer to two decimal places.

________%

d) What is the firm’s weighted-average cost of capital? Round your answer to two decimal places.

______ %
Answer #1

a) after tax cost of debt = 7% * 0.70 = 4.90%

b) cost of preferred stock = 10/96 = 10.42%

c) cost of common stock = 0.90*1.03/15 + 3% = 9.18%

d

Amount weight cost weight*cost
equity                   292,500.00 0.6500 9.180% 0.0597
debt                   112,500.00 0.2500 4.900% 0.0123
pref stock                     45,000.00 0.1000 10.420% 0.0104

WACC = 8.23%